# REFINER: Reasoning Feedback on Intermediate Representations

### v1

@@ 4 Experimental Setup
Datasets.
We evaluate REFINER on three diverse tasks (examples in Fig. 2). We briefly describe the datasets used for each task below [^7] In Table 6, we report the data statistics in Appendix.
Math Word Problem (MWP): We train our models on MAWPs (Koncel-Kedziorski et al., 2016) dataset and evaluated our models on a challenging dataset SVAMP (Patel et al., 2021). We evaluate our model on both the equation generation ($z$) and answer prediction ($y$) tasks. Similar to Ling et al. (2017); Amini et al. (2019) for equation generation, we replace the numeric values with variable names, for example, number0, number1, etc. For Synthetic Natural Language Reasoning (sNLR), we use the dataset from Liang et al. (2022) with the difficulty level as hard. We evaluate our model on both inference rule generation ($z$) and consequent generation ($y$). For Moral Story (MS), we use dataset from (Emelin et al., 2021), where we evaluate our model on moral norm z{z} and the moral action y{y} generation.
Training Details.
For each task, we train a UnifiedQa-T5-base model (UQA-base) (Khashabi et al., 2020) as a critic (§3.1). For exploration (§3.2), we use nucleus sampling with $p=0.5$. We select the hyper-parameters by the validation loss: for both the generator and critic model, we use the Adam optimizer with a learning rate of $1e^{-4}$. Each model is trained for $20$ epochs with early stopping based on validation loss. We trained all models on one A100 GPU. We run our models with $3$ random seeds and report the average results. For the human study, we selected outputs from the best models (baselines and our model) according to automatic metrics. We train models with $T=3$ iterations.
At inference time, we use greedy decoding for the generator and critic model with $T=1$ for the automatic critic and $T=3$ for the oracle critic. On the MWP and sNLR tasks, we use the exact match (EM) metric for intermediate steps (equation generation and inference rules) and accuracy (Acc) for the final answers. For MS, we conduct a manual evaluation study to assess the relevance of norms and moral actions [^8] Since the automatic scores such as BLUE, ROUGE, etc. only account for word level similarity between gold norms or actions and generate norms or actions.. Further evaluation details are provided in Appendix F.
Baselines.
We compare our method with three different LMs as generator models: UQA-base, UQA-large (supervised setting), GPT-$3.5$-code-DaVinci-002 (few-shot setting) and GPT-$3.5^{\star}$-text-DaVinci-003. We also compare REFINER to Proximal Policy Optimization (PPO) RL-based method (Schulman et al., 2017). We use the implementation of PPO from (Ramamurthy et al., 2022). For GPT-3.5, the model is not finetuned but prompted with standard few-shot prompts (Brown et al., 2020), in which the LM is given in-context demonstrations of input-output pairs. We provide $2$ for demonstrations per class. We also experimented with chain of thought (COT) prompting (Wei et al., 2022) where the model is prompted first to generate the intermediate steps ($z$) and then the final answer ($y$). Note that sNLR task is a synthetic task where the model needs to perform either one-hop or two-hop reasoning. Clark et al. (2021) showed that fine-tuning large language models (354354M parameter size) could achieve (99% accuracy) high performance. Hence, we only compare our REFINER model with the UQA-base model (220220M) (see Table 3). Since human annotation is expensive, we focus on comparing against the most meaningful baseline: UQA-large for MS task (see Table 4). It is important to highlight that our proposed framework is general and one can use any other LMs as generator or critic.
5 Results
Model Eq. ($z$) Ans. (y{y})
@@ 5 Results
REFINER large + Oracle (T=3) 68.1
GPT-$3.5$ + CoT 59.3 63.5
GPT-$3.5$ + CoT + REFINERcritic 62.3 66.4
GPT-$3.5^{\star}$ + CoT 64.1 67.1
GPT-$3.5^{\star}$ + CoT + REFINERcritic 67.3 70.6
@@
Performance on Intermediate Steps.
Table 2 reports the performance on the MWP task. We explored two different scenarios: (i) where the model only generates the equations ($z$) with variable names replacing the numeric values, and (ii) where the model generates both the equations and the final answers together. We observe for both scenarios that REFINER significantly outperforms baseline models with comparable sizes. Notably, UQA-base benefits (+13.1+13.1 EM) most when adding a critic in the loop. We observe that GPT-$3.5$ significantly benefits from REFINER trained critic. Since LLMs like GPT-$3.5$ (175175B parameters) are expensive to finetune, the improvement in equation generation of +3+3 and +3.2+3.2 EM without any modification is important. Interestingly, we observe that GPT-$3.5$ + COT manages to have significantly higher accuracy in answer $y$ than in equation $z$ (see Table 2). This result is similar to the observation made by Ye and Durrett (2022) and suggests that the intermediate equations can be unreliable. Also, we observe that if the critic was perfect ( Oracle), then REFINER could perform very strongly. Finally, REFINER could even outperform PPO, which uses BLEU-score as a reward function. This suggests that semi-structured fine-grained textual feedback can be more beneficial than value-based (where values are from automatic metrics) reward feedback. Note that this result may vary when these models are optimized directly with complex human values, as shown in Stiennon et al. (2020). Qualitatively, REFINER can correct incorrect equations through structured feedback, fixing the operators within a multistep solution (see Figure 8).
Model IR ($z$) Con (y{y})
@@
REFINER base + Oracle 96.2 $\pm$ 0.9 98.9
GPT-$3.5$ + CoT 17.4 $\pm$ 0.5 45.0
GPT-$3.5$ + CoT + REFINER 26.8 $\pm$ 0.5 46.6
GPT-$3.5^{\star}$ + CoT 14.3 $\pm$ 0.9 40.6
GPT-$3.5^{\star}$ + CoT + REFINER 21.1 $\pm$ 1.2 42.1
Table 3: Results on sNLR task. IR: Inference Rules, Con: Consequent
Model Norm ($z$) Action ($y$)
@@
For sNLR, similar to Liang et al. (2022), we observe that GPT-3.5 (code & text) performs poorly (see Table 3). We also find that GPT-3.5 code-DaVinci is better than text-DaVinci on the sNLR task. REFINER improves +2.9+2.9, +6.8+6.8 and +9.4+9.4 EM scores over UQA-base, GPT-$3.5^{\star}$ and GPT-$3.5$, respectively. Contrary to the MWP, the final answer $y$ is not a symbolic execution away from the intermediate step $z$, but we still observe that REFINER focuses on improving the intermediate step $z$, resulting in significant improvements in the answer $y$ prediction. Again, we observe that REFINER with a UQA-base can outperform few-shot prompted GPT3.5. Thus, our critic can identify the fine-grained reasoning errors and help improve the performance on inference rules generation.
@@
For MS, we assess the generation quality with three human judges who indicate whether the generated norms and moral actions are relevant to the given moral story. In Table 4, we summarize human evaluation results on $100$ moral story examples randomly sampled from the MS test dataset. More specifically, we report evaluation breakdown for both norm and moral action by the number of instances that are either Irrelevant, Unsure or Relevant along with Krippendorf’s $\alpha$ Krippendorff (2018) agreement scores. The results show an improvement of $20$ points, increasing the relevance over a strong UQA-large baseline. Hence, this suggests that feedback from a critic model, which has $3$ times fewer parameters than the generator, can improve the performance on generating reasoning steps. Further evaluation details are provided in Appendix F.
Performance on Final Answer Prediction.
@@
We observe that REFINER outperforms the strong LM baselines by +2.9,+3.2,+15+2.9,+3.2,+15 points for MWP, sNLR, and MS, respectively. These results support our hypothesis that improving the performance of generating better intermediate steps can result in better answer prediction. Notably, on the sNLR task, for GPT-$3.5$, we observe that by adding a critic, there is an improvement of $9.4$ in inference step generation; however, only +1.6+1.6 is in consequent prediction. This result indicates that for LLMs, these intermediate inferences can be inappropriate for the actual computation leading to the model’s answer. Hence, training to improve the quality of the intermediate step can result in better performance on the final prediction.
@@
REFINER base 47.2
@@
Table 5: Ablation Result on MWP task; Comparing model without critic during inference, and without the exploration (exp) phase during training. The results are measure with exact match of the generated equation, comparable to Table 2.
Ablation
To obtain better insight into the contributions of the individual components of our models, we perform an ablation study (Table 5). We observe that there is a considerable drop in performance from $47.2$ to $39.8$ when we do not use the critic model during inference. Hence, this result indicates that our generator model can leverage the feedback from the critic at inference time. Further, we find that the exploration step improves the performance +3.3+3.3 over the baseline model. This result supports our hypothesis that the exploration step increases the output variance and gives the generator model the opportunity to learn over a wide range of feedback.
@@ 6 Analysis
Figure 4: Examples. REFINER on MWP task. There different scenarios are highlighted in the figure, where (a) the critic model provides a correct feedback, generator model utilizes the feedback and fixes the incorrect equation, (b) the critic model provides a correct feedback however, generator model fails to fix the incorrect equation, and (c) the critic model provides an incomplete feedback generator model partially fixes the incorrect equation.
Figure 5: Error analysis. Number of errors made by baseline UQA-large and REFINER on 100 instances sampled randomly from dev sets of both datasets. Errors are categorized according to Table 1).
@@
6.1 Quantitative Analysis
Error Analysis.
In order to get more insight into the performance of our method, we conduct a fine-grained error analysis on the MWP and MS datasets (Fig. 5). We note that the most frequent errors are Incorrect Numbers for MWP and Semantic Misalignment for MS. An intuitive reason can be that for the MWP task, the models are sensitive to the numbers order as argued in (Patel et al., 2021). For MS, generating norms grounded in the context is challenging. Our analyses show a clear trend that REFINER is able to considerably reduce the errors for both datasets. This indicates that our trained critic model could identify fine-grained reasoning errors during inference.
Noise Sensitivity.
To further understand the behaviour of the REFINER framework, we run variations with noisy critics. We replace the oracle critic used during training with a noisy critic in (Fig. 6 (a)) to inspect how training with an imperfect critic impacts the generator. Also, we use a noisy critic at inference but keep the oracle critic during training (in Fig. 6 (b)). This analysis is performed on the SVAMP dataset for MWP. The noisy critics are generated by random perturbations of the oracle critic; for a noise-level $\epsilon$, the oracle feedback is replaced by random feedback with probability $\epsilon$.
We observe, in Fig. 6 (a), that when training with a very noisy critic ($75\%$ noise), the generator LM learns to ignore the critic, as there is no difference between using the trained critic or the oracle during inference. Interestingly, training with a bit of noise ($50\%$) does not seem to harm the model, as performances are not statistically different than training with the oracle (noise of $0\%$). Then, in Fig. 6 (b), we remark that the quality of the critic used at inference time has a huge impact. Having oracle provide feedback is by far the best scenario. Already with $25\%$ noise, the critic makes the generator perform worse than using our trained critic (REFINER). With more than $50\%$ noise, the critic significantly harms the generator. Indeed, the generator, trained with an oracle critic, has learned to trust the critic and expects useful feedback.
6.2 Qualitative Analysis
To explain the findings in §6.1, we further manually analyze 100 instances for the MWP task. We observe two different scenarios when REFINER failed to fix the outputs generated by generator model: (a) when the critic model provides a correct feedback; however, the generator model still generates incorrect equation, and (b) the critic model provides an incomplete or partially correct feedback. The former case indicates that either the generator model makes mistakes in following the instruction from the critic or the feedback from the critic can be ambiguous. For example, in Fig. 4, (b) we observe the case when the critic is correct, but the feedback could result in an incorrect equation. The latter case indicates that our trained critic model generates incorrect feedback, which can result in incorrect or partially correct equations. We also observe that our critic model failed to generate correct feedback when the generator model generates incorrect equations with multiple mistakes (see Fig. 4(c)).

### v2

@@ 4 Experimental Setup
Generator Model Eq. ($z$) Ans. (y{y})
@@
GPT-$3.5$ + CoT 64.1 67.1
GPT-$3.5$ + CoT + REFINERcritic 67.3 70.6
Table 2: Results on MWP. Comparison of REFINER with baselines on the SVAMP dataset. The average score over three runs is reported (p<0.05). For models other than GPT-3.5, the answer can be obtained via symbolic execution of the equation and is thus a function of the validity of the equation.
Datasets. We evaluate REFINER on three diverse tasks (examples in Fig. 6). We briefly describe the datasets used for each task below.Math Word Problem (MWP): We train our models on MAWPs (Koncel-Kedziorski et al., 2016) dataset and evaluated our models on a challenging dataset SVAMP (Patel et al., 2021). We evaluate our model on both the equation generation ($z$) and answer prediction ($y$) tasks. Similar to Ling et al. (2017); Amini et al. (2019) for equation generation, we replace the numeric values with variable names, for example, number0, number1, etc. Further, we also evaluated on GSM8K (Cobbe et al., 2021b) dataset which consists of 8.5K high-quality linguistically diverse grade school math word problems. For Synthetic Natural Language Reasoning (sNLR), we use the dataset from Liang et al. (2022) with the difficulty level as hard. We evaluate our model on both inference rule generation ($z$) and consequent generation ($y$). For Moral Story (MS), we use a dataset from (Emelin et al., 2021), where we evaluate our model on moral norm z{z} and the moral action y{y} generation.
Training Details. For each task, we train a UnifiedQa-T5-base model (UQA-base) (Khashabi et al., 2020) as a critic (§3.1). For exploration (§3.2), we use nucleus sampling with $p=0.5$. We select the hyper-parameters by the validation loss: for both the generator and critic model, we use the Adam optimizer with a learning rate of $1e^{-4}$. Each model is trained for $20$ epochs with early stopping based on validation loss. We trained all models on one A100 GPU. We run our models with $3$ random seeds and report the average results. For the human study, we selected outputs from the best models (baselines and our model) according to automatic metrics. We train models with $T=3$ iterations.
At inference time, we use greedy decoding for the generator and critic model with $T=1$ for the automatic critic and $T=3$ for the oracle critic. On the MWP and sNLR tasks, we use the exact match (EM) metric for intermediate steps (equation generation and inference rules) and accuracy (Acc) for the final answers. For MS, we conduct a manual evaluation study to assess the relevance of norms and moral actions [^6] Since the automatic scores such as BLUE, ROUGE, etc. only account for word level similarity between gold norms or actions and generate norms or actions.. Further evaluation details are provided in Appendix G. To train the critic model, we used the feedback data generated using the rule-based perturbation strategy (see §3.1).
Baselines. We compare our method with three different LMs as generator models: UQA-base, UQA-large (supervised setting), GPT-$3.5$-text-DaVinci-003 and ChatGPT (few-shot setting). We also compare REFINER to Proximal Policy Optimization (PPO) RL-based method (Schulman et al., 2017). We use the implementation of PPO from (Ramamurthy et al., 2022). For GPT-$3.5$, we provide $2$ for demonstrations per class. We also experimented with chain of thought (COT) prompting (Wei et al., 2022) where the model is prompted first to generate the intermediate steps ($z$) and then the final answer ($y$). Note that the sNLR task is a synthetic task where the model needs to perform either one-hop or two-hop reasoning. Clark et al. (2021) showed that fine-tuning large language models (354354M parameter size) could achieve (99% accuracy) high performance. Hence, we only compare our REFINER model with the UQA-base model (220220M) (see Table 3). Since human annotation is expensive, we focus on comparing against the most meaningful baseline: UQA-large for MS task (see Table 4). It is important to highlight that our proposed framework is general, and one can use any other LMs as generator or critic.
5 Results
@@
Generator Model IR ($z$) Con (y{y})
@@
GPT-$3.5$ + CoT 14.3 $\pm$ 0.9 40.6
GPT-$3.5$ + CoT + REFINER 21.1 $\pm$ 1.2 42.1
Table 3: Results on sNLR task. The average score over three runs is reported (p<0.05). IR: Inference Rules (Exact Match), Con: Consequent (Accuracy)
Performance on Intermediate Steps. Table 2 reports the performance of the MWP task. We explored two different scenarios: (i) where the model only generates the equations ($z$) with variable names replacing the numeric values, and (ii) where the model generates both the equations and the final answers together. We observe for both scenarios that REFINER significantly outperforms baseline models with comparable sizes. Notably, UQA-base benefits most (+13.1+13.1 EM) when adding a critic in the loop. We observe that GPT-$3.5$ significantly benefits from the REFINER trained critic. Since LLMs like GPT-$3.5$ (175175B parameters) are expensive to finetune, the improvement in equation generation of +3.2+3.2 EM without any modification is important. Interestingly, we observe that GPT-$3.5$ + COT manages to have significantly higher accuracy in answer $y$ than in equation $z$ (see Table 2). This result is similar to the observation made by Ye and Durrett (2022) and suggests that the intermediate equations can be unreliable. Finally, REFINER could even outperform PPO, which uses BLEU-score as a reward function. This suggests that semi-structured fine-grained textual feedback is more beneficial than value-based (where values are from automatic metrics) reward feedback. Note that this result may vary when these models are optimized directly with complex human values, as shown in Stiennon et al. (2020). Qualitatively, REFINER can correct incorrect equations through structured feedback, fixing the operators within a multistep solution (see Fig. 7).
Norm ($z$) Action ($y$)
@@
For sNLR, similar to Liang et al. (2022), we observe that GPT-3.5 performs poorly (see Table 3). REFINER improves +2.9+2.9, and +6.8+6.8 EM scores over UQA-base, and GPT-$3.5$, respectively. Contrary to the MWP, the final answer $y$ is not a symbolic execution away from the intermediate step $z$, but we still observe that REFINER focuses on improving the intermediate step $z$, resulting in significant improvements in the answer $y$ prediction. Again, we observe that REFINER with a UQA-base can outperform few-shot prompted GPT-$3.5$. Thus, our critic can identify the fine-grained reasoning errors and help improve the performance on inference rules generation.
@@
For MS, we assess the generation quality with three human judges who indicate whether the generated norms and moral actions are relevant to the given moral story. Table 4 summarises human evaluation results on $100$ moral story examples randomly sampled from the MS test dataset. More specifically, we report evaluation breakdown for both norm and moral action by the number of instances that are either Irrelevant, Unsure or Relevant along with Krippendorf’s $\alpha$ Krippendorff (2018) agreement scores. The results show an improvement of $20$ points, increasing the relevance over a strong UQA-large baseline. Hence, this suggests that a specialized critic model with $3$ times fewer parameters than the generator can improve the performance on generating reasoning steps.
@@
Performance on Final Answer Prediction. We observe that REFINER outperforms the strong LM baselines by +3.5,+3.2,+15+3.5,+3.2,+15 points for MWP, sNLR, and MS, respectively. These results support our hypothesis that generating better intermediate steps can result in better answer prediction. Notably, on the sNLR task, for GPT-$3.5$, we observe that by adding a critic, there is an improvement of +$6.8$ in inference step generation; however, only +1.5+1.5 in the consequent prediction. This result indicates that LLMs may either not use these intermediate steps to perform the deduction or fail to perform deduction.
Generator Model SVAMP GSM8K
GPT-3.5 ChatGPT GPT-3.5 ChatGPT
CoT 67.1 68.2 63.5 74.1
Self-reflection 67.2 68.4 63.1 74.6
Self-refine 67.6 68.2 63.8 74.7
REFINER 70.6 71.4 66.2 75.9
ReACT 67.3 68.4 64.7 75.5
ReACT + REFINER 70.6 71.9 67.8 77.4
Self-consistency 69.5 70.4 65.5 76.1
Self-consistency + REFINER 72.1 72.5 67.2 78.1
Table 5: Comparison with different refinement methods on SVAMP and GSM8K datasets. Averaged accuracy over three runs on the test sets is reported (p<0.05).
Comparing REFINER with other refinement methods. In Table 5, we compare REFINER with two other recent refinement methods: Self-refine (Madaan et al., 2023) and Self-reflection (Shinn et al., 2023) method on the SVAMP and GSM8K datasets. Both these baseline methods use LLMs to generate automatic feedback. Similar to Madaan et al. (2023), we observe that self-refine has minor improvement for MWP tasks. On the contrary, we find that REFINER significantly improves the performance of GPT-3.5 and ChatGPT by +$3.3$ and +$2.2$ on SVAMP and GSM8K datasets, respectively. This highlights the benefit of training a specialised critic that is grounded to the task. It can make LLMs more accurate than feedback from a general-purpose model (GPT-$3.5$ or ChatGPT). In Appendix §6, we have provided more details about the quality of feedback generated using our trained critic and GPT-$3.5$ (see Table 8). Further, we assess the performance of REFINER in improving the CoT generated by two recent methods: Self-Consistency (Wang et al., 2023) and ReACT method (Yao et al., 2023). We observe that REFINER can improve self-consistency and ReACT by +$2.02$ and +$2.9$. This demonstrates that a trained critic can be used as a tool and can bring performance gains to different methods out-of-the-box (more details in Appendix §A.2).
@@
REFINER base + critic datarule-based 47.2
@@
REFINER base + critic datasynthetic 44.1
REFINER base + criticOracle 66.0
Table 6: Ablation Result on MWP task; Comparing model without critic during inference, and without the exploration (exp) phase during training. We report the exact match scores of the generated equation, comparable to Table 2.
Ablation. To obtain better insight into the contributions of the individual components of our models, we perform an ablation study (Table 6). We observe that there is a considerable drop in performance from $47.2$ to $39.8$ when we do not use the critic model during inference. Hence, this result indicates that our generator model can leverage the feedback from the critic at inference time. Further, we find that the exploration step improves the performance +3.3+3.3 over the baseline model. This result supports our hypothesis that the exploration step increases the output variance and gives the generator model the opportunity to learn over a wide range of feedback. We compared the performance with the critic model trained on two different training data (see §3.1). We find that the critic trained on small automatically generated data using GPT-3.5 works better than without the critic in the loop. This result motivates researchers to use this method to generate negative samples to train their critic or preference learning model. Finally, we also observe that if the critic was perfect (Oracle), then REFINER can significantly improve the performance by fixing the mistakes generated by the generator model. This result indicates that REFINER can be seen as a framework that allows AI-AI and human-AI interaction.
@@ 6 Analysis
Figure 3: Error analysis. Number of errors made by baseline UQA-large and REFINER on 100 instances sampled randomly from test sets of both datasets. Errors are categorized according to Table 1).
Error Analysis. In order to get more insight into the performance of our method, we conduct a fine-grained error analysis on the MWP and MS datasets (Fig. 3). We note that the most frequent errors are Incorrect Numbers for MWP and Semantic Misalignment for MS. An intuitive reason can be that for the MWP task, the models are sensitive to the numbers order as argued in (Patel et al., 2021). For MS, generating norms grounded in the context is challenging. Our analyses show a clear trend that REFINER is able to considerably reduce the errors for both datasets. This indicates that our trained critic model could identify fine-grained reasoning errors during inference.
Noise Sensitivity. To further understand the behaviour of the REFINER framework, we run variations with noisy critics for the MWP task. We replace the oracle critic used during training with a noisy critic in (Fig. 4 (a)) to inspect how training with an imperfect critic impacts the generator. We also use a noisy critic at inference while keep the oracle critic during training (in Fig. 4 (b)). The noisy critics are generated by random perturbations of the oracle critic; for a noise-level $\epsilon$, the oracle feedback is replaced by random feedback with probability $\epsilon$.
Fig. 4 (a) shows that when training with a very noisy critic ($75\%$ noise), the generator LM learns to ignore the critic, as there is no difference between using the trained critic or the oracle during inference. Interestingly, training with a bit of noise ($50\%$) does not seem to harm the model, as performances are not statistically different than training with the oracle (noise of $0\%$). Fig. 4 (b) depicts the quality of the critic used at inference time has a huge impact. Having oracle provide feedback is by far the best scenario. Already with $25\%$ noise, the critic makes the generator perform worse than using our trained critic (REFINER). With more than $50\%$ noise, the critic significantly harms the generator. The generator, trained with an oracle critic, has learned to trust the critic and expects useful feedback.
@@
Qualitative Analysis. To explain the findings in §6, we further manually analyze 100 instances for the MWP task. We observe two different scenarios when REFINER failed to fix the outputs generated by generator model: (a) when the critic model provides a correct feedback; however, the generator model still generates incorrect equation, and (b) the critic model provides an incomplete or partially correct feedback. The former case indicates that either the generator model makes mistakes in following the instruction from the critic or the feedback from the critic can be ambiguous. For example, in Appx Fig. 5, (b) we observe the case when the critic is correct, but the feedback could result in an incorrect equation. The latter case indicates that our trained critic model generates incorrect feedback, which can result in incorrect or partially correct equations. We also observe that our critic model failed to generate correct feedback when the generator model generates incorrect equations with multiple mistakes.
Task UQA (220M) UQA (770M) GPT-3 (175B)
MWP 69.5 +/- 2.6 73.4 +/- 3.7 63.5 +/- 5.6
sNLR 95.5 +/- 1.4 98 +/- 2.2 34.5 +/- 2.4
MN 77.4 +/-2.5 80 +/- 4.5 76.4 +/-3.5
Table 7: Comparing the performance of different critic models. Exact-match score is reported.
Quality of the feedback. To better understand the difference in the quality of the feedback, we compare our trained critic model with GPT-3.5. We assess the quality of the feedback on 500 instances per task and report the exact match scores in Table 8. Please note that we include instances where the critic feedback should say the solution is correct and hence generate ’No’. For GPT-3.5, we have provided (two) few-shot examples per type of error and two examples with ’No’ as feedback. Our results show that trained critic (UQA) can comprehensively outperform GPT-3.5. We observe that GPT-3.5 performs well in identifying when the answer is correct. However, it makes errors when asked to generate meaningful semi-structured feedback for incorrect reasoning steps.